I need help solving this problem.
A car travels 300 km in the same amount of time that a freight train travels 200 km. The speed of the car is 20km/hr more than the speed of the train. Find the speed of both the car and the train.

Answer by mananth(16949)

let speed of train be x km/hr
speed of car = x+20 km/hr
..
distance by car = 300 km
distance by train = 200 km
..
Time taken by both is same
200/x = 300 (x+20)
300x = 200(x+20)
300x = 200x +4000
add -200x
300x-200x=4000
100x = 4000
x = 40 km/hr speed of train
x+20 = 40 +20 = 60km/hr. speed of car